### v2

# Are We Solving a Well-Defined Problem? A Task-Centric Perspective on Recommendation Tasks

Abstract.
Recommender systems (RecSys) leverage user interaction history to predict and suggest relevant items, shaping user experiences across various domains. While many studies adopt a general problem definition, i.e., to recommend preferred items to users based on past interactions, such abstraction often lacks the domain-specific nuances necessary for practical deployment. However, models are frequently evaluated using datasets from online recommender platforms, which inherently reflect these specificities. In this paper, we analyze RecSys task formulations, emphasizing key components such as input-output structures, temporal dynamics, and candidate item selection. All these factors directly impact offline evaluation. We further examine the complexities of user-item interactions, including decision-making costs, multi-step engagements, and unobservable interactions, which may influence model design and loss functions. Additionally, we explore the balance between task specificity and model generalizability, highlighting how well-defined task formulations serve as the foundation for robust evaluation and effective solution development. By clarifying task definitions and their implications, this work provides a structured perspective on RecSys research. The goal is to help researchers better navigate the field, particularly in understanding specificities of the RecSys tasks and ensuring fair and meaningful evaluations.
Recommender Systems, Task Formulation, Evaluation, User cost
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Draft; March 2025; ACM Templateisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
In many recommender systems (RecSys) studies, solutions are proposed for a commonly adopted problem: given a set of users, items, and their interactions, the goal is to recommend items that align with users’ interests or preferences. While this general problem definition captures common patterns across recommendation scenarios, its abstraction overlooks critical details needed for practical RecSys applications. Furthermore, discussions on task definition often lack clarity, particularly regarding its scope and practical implications. At the same time, RecSys research is closely tied to real-world applications, where models are primarily evaluated using datasets obtained from operational recommender platforms. The mismatch between abstract problem definitions and domain-specific evaluations leads to inconsistent settings and findings across experiments (Zangerle and Bauer, 2022; McElfresh et al., 2022). In this paper, we present a comprehensive analysis of different RecSys tasks, highlighting their key distinctions and implications for evaluation and real-world deployment.
Recommender systems are not a new research area, and numerous influential works exist. We begin our discussion in Section 2 with a review of several highly cited works in RecSys, emphasizing task formulations. Interestingly, key factors in RecSys were well defined and discussed decades ago. Yet, the community still disagrees on the choice of baselines and datasets (Ivanova et al., 2023; Beel et al., 2024).
One main contribution of this paper is a detailed examination of task definition in Section 3, focusing primarily on the input and output of a mapping function, i.e., a recommender. We discuss a missing element in task formulation i.e., time, and the selection of candidate items for recommendation, both of which have a significant impact on offline evaluation.
Another contribution of this paper is a detailed analysis of the life cycle of user-item interactions in Section 4. In the pre-interaction judgment stage, users are presented with a list of choices or recommended items. Users incur costs when making informed (or sometimes uninformed) decisions to select a recommendation, which serves as a clear indication of user preferences. We then explore the complexity of user-item interactions. For example, some interactions can involve multiple steps, as in e-commerce, where a user may add an item to their cart, proceed to payment, await delivery, interact with the product, and possibly return it. In music streaming, users are not offered a choice but can only skip or fastforward a recommended song. In some cases, true user-item interactions may not be directly observable as in job recommendation.
Based on the task definition and the framework for understanding it, we present our perspectives in Section 5. We consider (i) the ultimate goal of recommendation, i.e., reducing user cost, (ii) the relationship between task formulation, solution, and evaluation, and (iii) the intersection of task specificity and model generalizability. In our discussion, we use a few examples to show the differences in RecSys tasks based on the presented dimensions, i.e., candidate item selection, pre-interaction judgment, and user-item interaction. We hope that this paper provides researchers, particularly those new to recommender systems, with a clearer understanding of the research topic. After establishing a clear understanding, subsequent research can precisely define specific RecSys tasks, select appropriate baselines, and design experiments that account for task specificity.
2. A Historical Review of Task Formulation
We begin with four of the most widely cited papers, [^1] Citation counts were obtained via a Google Scholar search for papers with the title words “recommender systems” or “recommender system” on February 12, 2025. each with over 8,000 citations. As foundational works in this field, these papers have influenced many researchers, and the tasks they define have likely shaped numerous follow-up studies. The four papers reviewed were published between 2004 and 2011 (Herlocker et al., 2004; Koren et al., 2009; Ricci et al., 2011; Adomavicius and Tuzhilin, 2005).
The Recommendation Problem. In one of the most influential survey papers in RecSys, Adomavicius and Tuzhilin (2005) formally define the recommendation problem as follows:
Definition 0 (Recommendation Problem).
Let $U$ be the set of all users and let $I$ be the set of all possible items that can be recommended. Let $r$ be a utility function that measures the usefulness of item $i$ to user $u$, i.e., $r\in U\times I\rightarrow R$, where $R$ is a totally ordered set (i.e., nonnegative integers or real numbers within a certain range). Then, for each user $u\in U$, we want to choose such item $i^{\prime}\in I$ that maximizes the user’s utility:
(1) $\forall u\in U,i^{\prime}_{u}=\arg\max_{i\in I}r(u,i)$
The authors further note that “the utility of an item is usually represented by a rating, which indicates how a particular user liked a particular item,” a concept commonly known as explicit feedback. Nevertheless, in most subsequent RecSys studies, implicit feedback has become far more prevalent (Rendle et al., 2009). As a result, the utility of an item to a user is often inferred from binary feedback, whether the user has interacted with the item, such as listening to a song, purchasing a watch, or staying at a hotel. Notably, the difference between explicit and implicit feedback primarily affects data modeling and the loss function design when learning $r$, while the core recommendation problem remains unchanged.
Collaborative Filtering. Koren et al. (2009) do not formally define tasks in recommender systems but instead compare the two primary approaches: content filtering and collaborative filtering. They describe content filtering as a method that creates user or item profiles to capture their characteristics, enabling the recommender to associate users with relevant items. In contrast, collaborative filtering relies exclusively on past user-item interactions, such as previous transactions or product ratings. The authors highlight that “a major appeal of collaborative filtering is that it is domain-free, yet it can address data aspects that are often elusive and difficult to profile using content filtering,” further noting that collaborative filtering is “generally more accurate than content-based techniques.”
Due to these advantages, collaborative filtering solutions are often considered more generalizable, aligning well with one of the key objectives in academic research. However, the authors also acknowledge the cold start problem in collaborative filtering, which arises when new users or items have no past interactions (Schein et al., 2002).
Although solutions such as content-based, collaborative, and hybrid filtering are independent of problem definitions, the widespread adoption of collaborative filtering in recent research leads us to assume that user-item interactions remain a key input for typical recommender systems. With this in mind, we reiterate the points made in (Koren et al., 2009), emphasizing that collaborative filtering relies exclusively on past user-item interactions, such as previous transactions or product ratings.
User Feedback and Preference. In the introduction chapter of the RecSys Handbook, Ricci et al. (2011) describe personalized recommendations as ranked lists of items, where recommender systems aim to “predict what the most suitable products or services are, based on the user’s preferences and constraints.” Regarding user preferences, the authors highlight that these can also be inferred from their actions, such as navigating to a specific product page. In this context, user feedback can take multiple forms: explicit feedback, such as ratings; implicit feedback, derived from user-item interactions; and inferred preferences based on observed user behavior.
Task Diversity. Herlocker et al. (2004) provide a thorough discussion on user tasks for recommender systems. Their discussion focuses on end-user tasks (i.e., not marketers or other system stakeholders), [^2] We refer readers to (Zangerle and Bauer, 2022) for more detailed discussion on other users in a recommender system like item provider, platform provider, and other stakeholders. which aligns well with the RecSys tasks to be discussed in this paper. The key user task is to find good items, such as providing users with a ranked list of recommended items. The authors highlight that “there are likely to be many specializations of the tasks within each domain,” and the domain-specific characteristics are reflected in the properties of the datasets, e.g., the need for novelty versus quality, as well as the cost/benefit ratio of false/true positives/negatives. Regarding novelty, in certain domains, “the user goal is dominated by finding recommendations for things a user does not already know about.” In other domains, users expect recommendations to help them find high-quality items among many choices. Regarding cost/benefit, the authors provide an example of the relatively low cost of false-positive movie recommendations. However, in other settings, e.g., if an item comes with a high price tag, then false positives may carry a much higher cost.
Summary. From this historical review, we derive the following main points: (i) The primary task of a recommender system is to suggest items to users based on their preferences. (ii) Collaborative filtering is a widely adopted approach that leverages past user-item interactions. (iii) In most cases, users provide implicit feedback, and their preferences can also be inferred from their actions. (iv) Recommendation tasks can be domain-specific, with various factors to consider, such as the desired level of novelty and quality, as well as the trade-off between cost and benefit.
3. A Closer Look at the Task Definition
Based on the above discussion and the main points identified, we now examine the RecSys task definition from multiple perspectives. For ease of presentation, we rewrite Definition 2.1 by specifying recommender as a mapping from its input to output.
Figure 1. The classic task definition of recommendation, where user-item interactions are represented in a matrix form. The objective is to predict the missing values, which correspond to unobserved user-item interactions.
Definition 0 (Recommendation).
Let $U$ be a set of users and $I$ be a set of items. A recommender system aims to produce a ranked list of items for a user $u$, based on user-item interactions $U\times I$.
(2) $\langle u,U\times I,I\rangle\rightarrow R^{u}$
In this rewritten definition, the recommender system’s input consists of three components: (i) the user $u$ for whom recommendations are to be made, (ii) the set of user-item interactions $U\times I$, which includes all existing interactions made by all users and all items, and (iii) the set of available items $I$ from which recommendations are being made. The system outputs a ranked list of recommended items for $u$, denoted by $R^{u}$.
Following the tradition of illustrating recommender system tasks in matrix form, we present Definition 3.1 in Figure 1. In this figure, the goal is to generate a ranked list of recommended items for $u_{3}$. The items $i_{1}$ to $i_{n}$ collectively form the item set $I$, and the interactions encompass all interactions made between the $m$ users and $n$ items.
Note that this definition is generic to cover all attributes or side information of users or items, because all such information or metadata can be easily derived given a user ID or item ID. As we focus on the core problem definition of RecSys in this paper, we omit all these details.
3.1. The Missing Input: Time
Recall that in (Koren et al., 2009), the authors note that collaborative filtering relies exclusively on past user-item interactions to make recommendations. In practical scenarios, whether recommending a product for purchase or a song to listen to, it is common sense that recommendations made at a time point $t$ should be based on all information available at or before $t$. Ideally, the user will find the recommendations appealing, then choose to interact with them after the new recommendations are made at $t$.
Adding Time to the Input. With that, we rewrite the mapping function in Definition 3.1 to consider the time dimension. We also make an assumption that each interaction between a user $u$ and item $i$ is associated with a time stamp $t_{x}$ when the interaction occurred, denoted by $(u,i,t_{x})\in U\times I$.
(3) $\langle u,t,(U\times I)_{\leq t},I_{\leq t}\rangle\rightarrow R_{t}^{u}$
In this rewritten form, we introduce a time point $t$, indicating a recommendation is to be made for user $u$ at time point $t$. We use $(U\times I)_{\leq t}$ to represent all user-item interactions that occurred before time $t$, a simplified form of {$(u,i,t_{x})\in U\times I|t_{x}\leq t$}. Accordingly, the candidate items available for recommendation are those that were registered in the system and accessible to the recommender as at $t$, denoted by $I_{\leq t}$.
Figure 2. The recommendation task along time line. Recommendations are to be made for user $u_{2}$ at time point $t_{e}$. A model is expected to learn from interactions occurred before $t_{e}$ i.e., $(U\times I)_{\leq t_{e}}$ and make recommendation among items available at $t_{e}$, denoted by $I_{\leq t_{e}}$.
While this reformulation may seem trivial and may not impact real-world or online RecSys recommendations, strictly adhering to this task definition poses significant challenges for offline evaluations (Sun, 2023).
Impact to Offline Evaluation. In offline evaluations, user-item interactions in a dataset must be partitioned into training and test sets. The former is used to train the RecSys model, while the latter is used to assess its performance. According to the task formulation in Equation 3, evaluating a model’s performance at time $t$ for a user $u$ requires that the model only access or learn from the interactions that occurred before $t$, i.e., $(U\times I)_{\leq t}$, and generate recommendations from $I_{\leq t}$, the items available at or before $t$.
Figure 2 provides an illustration where the items interacted with by each user are plotted to the right of the user along a global timeline. For example, user $u_{1}$ interacted with item $i_{1}$ at time point $t_{1}$. If we use the user-based leave-last-one-out data split to evaluate our model’s accuracy for user $u_{2}$, the last interaction of $u_{2}$ at time $t_{e}$ is masked as the test instance. The recommender can learn from all user-item interactions that occurred before $t_{e}$, i.e., $(U\times I)_{\leq t_{e}}$. The items available for recommendation to $u_{2}$ at $t_{e}$ are those that have been interacted with by any user before $t_{e}$. Notably, items $i_{7}$ and $i_{8}$ received their first interaction after $t_{e}$, the system has no interaction data for these items at $t_{e}$, and they should not be recommended to $u_{2}$. [^3] Note that our discussion assumes the recommendation model is based on collaborative filtering, to avoid the possible misunderstanding that items $i_{7}$ and $i_{8}$ could be recommended based on profile matching.
Finally, note that the user-based leave-last-one-out data split results in each user having their own $(U\times I)_{\leq t}$ and $I_{\leq t}$, since not all users have their last interaction at a common time. Without enforcing a global time point, many widely adopted data partitioning schemes, such as random splitting or user-based leave-last-one-out, lead to data leakage (Ji et al., 2023). This occurs when user-item interactions from the future (relative to time $t$ in Equation 3 or $t_{e}$ in Figure 2) are used to train the model and make recommendations.
In other words, when recommending items for a specific user, the model does not rely solely on that user’s interaction history but also leverages interactions from other users who may share similar preferences. The time point constrains all user-item interactions accessible to the system.
The incorporation of factor $t$ into the task formulation also directly impacts the implementation of certain baseline models. For instance, the widely used popularity baseline, which simply counts item occurrences in the training data, does not accurately reflect real-life item popularity. In practice, item popularity is typically associated with a reference time point and a specific time range, e.g., the top-selling books from last week or last month. By introducing a time point $t$ in problem definition, item popularity can be derived with reference to that specific time point (Ji et al., 2020, 2022).
3.2. The Missing Constraints on Candidate Items
Next, we focus on the available items that can be recommended to a user $u$ at time $t$. Herlocker et al. (2004) discuss recommender system tasks from the perspective of novelty and quality, assuming that users generally expect recommended items to be new and previously unconsumed. However, this assumption may not always hold, as a recommender system is ultimately designed to help users find preferred items more easily and quickly.
In many scenarios, users may prefer to interact with items they are already familiar with and confident in (Anderson et al., 2014). Groceries shopping (Ariannezhad et al., 2022; Katz et al., 2022), e-commerce (Wang et al., 2019; Quan et al., 2023), food delivery ordering (Li et al., 2024), and song listening (Tsukuda and Goto, 2020; Reiter-Haas et al., 2021) are a few examples where a user may prefer to have earlier consumed items to be recommended for easy selection.
Let $I^{u}_{t}$ be the items that $u$ has interacted with in the past at time point $t$. We use $I_{t}^{\bar{u}}$ to denote the remaining items available at $t$ that are new to $u$, then the recommendation in Equation 3 can be divided into two formulations for repeated consumption $R_{t}^{ur}$ and for exploration $R_{t}^{ue}$ recommendations, respectively.
(4) $\langle u,t,(U\times I)_{\leq t},I_{\leq t}^{u}$ $\rangle\rightarrow R_{t}^{ur}$
(5) $\langle u,t,(U\times I)_{\leq t},I_{\leq t}^{\bar{u}}$ $\rangle\rightarrow R_{t}^{ue}$
Note that, how to present $R_{t}^{ur}$ and $R_{t}^{ue}$ to users through a web or app interface, e.g., as two separate rankings or as a merged ranking of items from both recommendations, is orthogonal to our discussion in this paper.
Figure 3. The recommendation for repeated consumption for user $u_{2}$ at time $t_{e}$. The search space for repeated consumption is the two items $i_{2}$ and $i_{5}$. To recommend new items, the candidate items will be the remaining items in $I_{\leq t_{e}}\setminus\{i_{2},i_{5}\}$.
Impact to Offline Evaluation. Again, the partitioning of candidate items for recommendation $I$ into $I^{u}$ and $I^{\bar{u}}$ may seem trivial (here, we omit the subscript $t$ for clarity), but this partitioning leads to (i) significant differences in the search space for a recommender, because $|I^{u}|\ll|I^{\bar{u}}|$ and $|I^{\bar{u}}|\approx|I|$, and (ii) subsequently significant differences or impacts on evaluation.
It is important to note that if repeated consumption and exploration are not clearly distinguished, the evaluation results may be significantly biased toward repeated consumption. Because of the much smaller item search space for recommendation and a large portion of, if not all, items being users’ preferred items, the chances of making a good recommendation among the repeated items are much higher than picking an item from a much larger pool of items, especially those the user has no experience with. The recommendation accuracy could be easily dominated by the repeated items if repeated consumption is common like food ordering and groceries shopping. Hence, the separation of evaluations of repeated consumption and exploration better reflects the model’s accuracy; a detailed discussion and comparison is presented in  (Li et al., 2024) for food delivery recommendation. Additionally, due to the much smaller search space, recommending repeated items is significantly more efficient than recommending items for exploration.
Candidate Item Selection. The repeated consumption and exploration is just one example of candidate item selection. The candidate items suitable for recommendation can be directly specified by users. For example, if a user is planning a trip to a city, hotel recommendations would only be meaningful if they are located in that city and available during the user’s travel dates. Although this may seem trivial and makes the recommendation problem more similar to a search problem, the vast majority of recommendation work using the Yelp dataset does not consider users’ geographical constraints, to the best of our knowledge. That is, if a user resides in New York City and primarily visits restaurants within a relatively small region of NYC, recommending a restaurant from a different state would be far less practical for this user. On the other hand, the candidate items suitable for recommendation could also be derived based on user intent mining through search logs and other means. Note that the user intent mining here is different from the line of research on intent-aware recommender systems (Jannach and Zanker, 2024; Shehzad et al., 2025), which will be discussed in Section 5.2. One key difference is that, in our discussion on candidate item selection, the set of eligible items to be recommended is determined before running the recommender system, hence is part of the input to the problem definition and can be specified by users, illustrated in Figure 4.
3.3. Yet Another Rewritten Task Formulation
We can now combine Equations 4 and 5 into a more general formulation by introducing a selection condition on the candidate items, denoted as $s(I_{\leq t})$. Specifically, $s(\cdot)$ can represent selection criteria based on user interaction history e.g., for repeated consumption, or be specified by users with temporal, spatial, or other attribute-based conditions. However, by listing $s(I_{\leq t})$ as part of the input to the recommendation task, the candidate set of items is determined before a model ranks them.
Definition 0 (Recommendation Task).
Let $U$ be a set of users and $I$ be a set of items. A recommender system aims to produce a ranked list of items for a user $u$ at time $t$, based on user-item interactions $U\times I$ that is available at $t$, and the conditions specified on the candidate items.
(6) $\left\langle u,t,(U\times I)_{\leq t},s(I_{\leq t})\right\rangle\rightarrow R_{t}^{u}$
Now, we define the recommendation task as making a recommendation at time $t$ for user $u$ by referencing user-item interactions that occurred before $t$ and selecting suitable candidate items that satisfy the selection criteria $s(\cdot)$ to meet the inferred user preference. Following the example of repeat consumption and exploration, the final recommendation items presented to a user could be the results of running multiple recommendations made from multiple sets of candidate items conditioned on different multiple $s(\cdot)^{\prime}s$. Nevertheless, in terms of model evaluation, it is more meaningful to independently evaluate each model specific to candidate items on one selection function $s(\cdot)$.
4. From Recommendation to User Consumption
Figure 4. User costs after recommendations made at time $t$: from pre-interaction judgment, to interaction, then to an optional post-interaction feedback. Among the recommended items, $i_{2}$ satisfies selection criteria $S(\cdot)$.
So far, we have discussed the importance of explicitly incorporating the time factor into the problem definition and examined suitable candidate items for recommendation based on the context of the recommendation task. We consider these to be essential factors in the task definition due to their direct impact on model evaluation, which in turn directly impact the model design. Next, we take a closer look at user-item interactions $U\times I$ and analyze their meaning, as instances $(u,i,t_{x})$ are widely used to train recommender systems in academic research.
There are many reasons for a user to interact with an item, such as through an intentional search rather than a recommendation. Since we are discussing recommendations, we make a general assumption that if a user interacts with an item, it reflects some form of preference for that item. However, we aim to explore the various stages from a recommended item $i\in R_{t}^{u}$ to an interaction $(u,i,t_{x})$, and to what extent a model can learn from the interaction. This process can be considered as another form a mapping:
(7) $\langle u,R_{t}^{u}\rangle\rightarrow(u,i,t_{x})\text{\quad where \quad}i\in R_{t}^{u}$
4.1. User-Item Interaction: The Life Cycle
In the early stages of RecSys research, many studies focused on explicit user feedback, such as ratings e.g., a user giving a movie a 5 out of 5 rating (Herlocker et al., 2004). However, since explicit feedback is often difficult to obtain, implicit feedback has become the mainstream in academic research e.g., a user has watched a movie. Nevertheless, Ricci et al. (2011) suggest that user preference can also be inferred from actions such as navigating to a product page e.g., a user clicking on a movie for more details among a list of candidate movies. However, there is a lack of detailed discussion on to what extent such feedback contribute to accurately modeling user preference.
In our discussion, we divide user-item interactions into three stages: pre-interaction judgment, interaction, and post-interaction feedback, illustrated in Figure 4. While these stages may not apply to all recommendation scenarios, they provide a useful framework for understanding the relationship between user-item interactions and user preferences.
Consider a user searching for a hotel in a city she has never visited, for a conference. She is presented with a list of 10 hotels on a booking website, initiating the pre-interaction judgment stage. She clicks on one hotel for more details, a decision influenced by factors such as images, branding, location, price, or other readily available attributes. Next, she enters the deliberate evaluation phase, where she reads the hotel description, checks room facilities, and reviews user feedback before deciding to book. All these effort happens before the user stays at the hotel.
A few months later, she attends the conference and stays at the hotel. This marks the interaction stage, where she gains firsthand experience. Finally, after checking out, she provides feedback in the form of a rating and review, entering the post-interaction feedback stage. This feedback may consider various factors in a typical hotel review, such as location convenience, price, food, service quality, and hygiene.
Both quick judgments and deliberate evaluations in the pre-interaction stage come at a cost, the user’s time and effort in gathering relevant information. The outcomes of these efforts strongly indicate user preference, especially in similar future situations, e.g., planning another trip or booking a restaurant.
Post-interaction feedback, however, may not always accurately reflect user preference. While a review can highlight the primary reasons for choosing a hotel, representing genuine user preference, it may also describe aspects such as expectation gaps and booking effort, which are not necessarily preference indicators. A rating, influenced by these factors, may not always be a reliable measure of preference compared to the concrete fact that the user chose to stay at the hotel. In fact, in many practical scenarios, post-interaction feedback is often difficult to collect, as it requires users to put in additional effort to provide a rating or write a review. Hence, in our following discussion, we focus on the other two stages.
In general, recommendations are presented as a list of candidate items before any user-item interaction occurs. At this stage, the user relies solely on their judgment, which inherently reflects preference, as they would likely make the same choice in similar circumstances. However, most academic datasets only contain actual interactions, lacking context on the reasons behind them. This limitation makes it challenging to model user preference purely based on observed interactions. Thus, a deeper understanding of the user decision-making process is crucial for more effective preference modeling.
4.2. Complexity of Pre-Interaction Judgment
We now focus on the pre-interaction judgment, which often takes the form of quick judgment or deliberate evaluation before a choice is made. The complexity of this judgment from the user’s perspective may arise from several dimensions.
Informed vs Uninformed Decision. One major factor from the user’s perspective is whether they possess the knowledge to accurately judge an item before interacting with it. For familiar items like books, movies, or other products the user has prior experience with, they can make an informed decision based on available attributes or other relevant information. Take movies as an example, users may decide whether to watch a film based on the director, cast, genre, a brief synopsis, or simply the poster. However, if a user has never used a robot vacuum before, many of the terms in the product description may be unfamiliar to them. Even after reading the machine specifications and user reviews, they may struggle to discern the pros and cons of a specific model, with references to her own home layout and floor conditions.
These uninformed decisions may not necessarily indicate user preferences. It is also challenging for a recommendation platform to determine whether a user’s decision was based on prior knowledge or made without a full understanding of the item. This distinction is crucial, as recommendations based on uninformed interactions might not truly reflect user interests, potentially leading to less effective personalization. For example, after using a robot vacuum for some time, the user may realize that he should purchase another model with features better suited to his floor conditions.
Items in One vs. Multiple Types/Categories. Many studies in recommendation are conducted on datasets containing only one type of item, such as books, movies, news, or music. In these cases, there are common characteristics, such as genre, director, or artist, that users can rely on for pre-interaction judgment before actually interacting with a recommended item. However, in the context of online shopping, where products span thousands of categories, users apply different criteria and expectations when making judgments for different types of items, for informed decisions.
In such a diverse setting, user preferences learned from interactions across multiple item types may be shaped more by general associative patterns than by genuine preferences for specific products. However, distinguishing between personal preferences and co-purchasing patterns remains challenging, as these so-called preferences may be expressed for broader item categories rather than specific items. Nevertheless, we acknowledge that user preference is a complex concept (Kleinberg et al., 2022).
4.3. Recognition of User-Item Interaction
After the pre-interaction judgment phase, the user makes a decision and interacts with an item. In most datasets used for RecSys evaluation, user-item interactions are recorded in the form of $(u,i,t_{x})$. However, interactions may appear in different forms depending on the recommendation context.
Interaction Process Can Be Complicated. Taking online shopping as an example, the process does not end when a user clicks on a recommended item. After deciding on a product, the user might add the item to their cart, proceed to payment, receive the delivery, and ultimately complete the purchase. From the shopping platform’s perspective, this sequence of events marks a successful interaction. However, complications can arise if the user later decides to return the product due to reasons such as quality issues, unmet expectations, or simply a change of mind.
This raises an important question: should an unsuccessful purchase (i.e., one that results in a return) still be considered a valid user-item interaction for learning user preferences? On one hand, the initial decision to purchase the item indicates some level of interest or preference. On the other hand, the return suggests dissatisfaction or misalignment with the user’s expectations. If returns are not accounted for, the model might incorrectly reinforce recommendations for similar items, leading to suboptimal suggestions. Therefore, when incorporating interaction data into preference modeling, it is crucial to differentiate between successful and unsuccessful purchases and consider additional contextual signals, such as return rates, to better understand user preferences.
Absence of Pre-Interaction Judgment. Some of the user-item interactions happen without a pre-interaction judgment phase. This is especially common in scenarios like music streaming and short-video viewing, where users often do not actively select each item. Instead, they are presented with an initial set of options, and after selecting the first item, subsequent content is automatically fed to users by the recommendation engine, e.g., a playlist or streaming.
In such cases, user engagement signals, such as skipping, fast-forwarding, or continuing to watch/listen, play a crucial role in modeling preferences. Unlike traditional recommendation settings where users consciously evaluate and select items before interaction, here, user preferences are inferred more dynamically based on real-time behaviors, or the interaction itself. In fact, users are less likely to explicitly express “like” in a streaming setting but are more likely to express “dislike”. This distinction raises important challenges for recommendation models, as they must distinguish between passive exposure and active preference while adapting to continuously evolving user interests. In this case, the common form of user-item interaction $(u,i,t_{x})$ becomes less accurate compared to other settings. Such recommendation in streaming form also post questions in item attribute modeling e.g., evaluating whether the cover image of a short video influence user viewing as the user may not even has the chance of viewing the cover image for each video in the streaming.
Unobservable Interaction. There are also recommendation scenarios where user-item interactions are not directly observable and can only be inferred from external sources.
One example is job recommendation, where matching is based on a user’s skills and knowledge against job requirements. Unlike traditional recommender systems where user engagement signals (such as clicks, purchases, or views) are readily available, the job application process involves significant effort on both ends, applicants must prepare resumes and cover letters, while companies conduct interviews before making offers. As a result, direct interactions, such as applying for a job or receiving an offer, may not always be captured by a job recommendation platform. Instead, implicit signals, such as a user frequently viewing job postings in a particular field or updating their profile, might be used to infer their interests and preferences.
This lack of direct interaction data introduces challenges in preference modeling, as user engagement may not always reflect strong interest, and external factors e.g., hiring decisions, can influence the outcome. Thus, in such cases, recommender systems must rely on richer contextual information and alternative feedback mechanisms to refine their predictions.
4.4. Interdependency Across Recommendations
Recommendations can occur either independently, as in hotel booking, or within a session-based context, such as music streaming and short-video viewing. In the latter case, subsequent recommendations may be influenced by the previous selections or even the initial choice made at the start of the session.
Figure 5. Reasons for a bad recommendation. The option “unrelated to the current video” suggests the dependency between the next recommendation and the current item.
Figure 5 shows a feedback form for assessing video recommendation quality from YouTube. Notably, one of the options for a poor recommendation is labeled as “unrelated to the current video.” This indicates that the recommendation system tailors the next suggested video based on the one currently being played. The inclusion of this feedback option emphasizes users’ expectations that recommendations within a session should maintain contextual relevance. This is a key characteristic of session-based recommendation, a specialized type of recommendation task (Li et al., 2025). A detailed discussion on the recommendation flow is made in (Sun, 2024).
This example further highlights the complexity of recommendation systems and the variety of factors that need to be accounted for when designing different recommendation tasks. Each recommendation scenario, whether based on user preferences, session context, or product types, requires careful formulation to ensure that the user experience is optimized and that the system accurately reflects user intent.
5. Discussion and Perspectives
Next, we discuss the ultimate goal of recommendation, i.e., reducing user cost, the relationship between task formulation, solution, and evaluation, and the intersection of task specificity and model generalizability, along with our perspectives on RecSys.
5.1. Recommendation vs User Cost
The ultimate goal of a recommender system is to reduce user effort in finding products or services of interest, enhance their enjoyment of recommendations, and build trust in the system. However, users still incur different costs at various stages of the interaction process. In the pre-interaction judgment stage, users must evaluate recommendations based on available information, such as descriptions, images, reviews, or other metadata, which requires cognitive effort and time. This cost is higher when users lack prior knowledge of the item category or when the recommendation list is overwhelming. During the interaction stage, users experience the actual product or service, which may involve monetary costs (e.g., purchasing a product), time investment (e.g., watching a recommended movie), or engagement effort (e.g., exploring an unfamiliar interface). If the recommendation is poor, users may feel frustrated, leading to dissatisfaction and disengagement. Notably, such negative experiences can accumulate over time, eventually causing users to leave the platform, confirmed by a recent large scale survey (Zou et al., 2024). Finally, in the post-interaction stage, users may be asked to provide feedback, such as ratings or reviews. While valuable for possibly improving recommendation quality, this step requires additional effort, and many users may choose not to participate.
Understanding and minimizing all such costs at each of the three stages is essential for enhancing user experience and optimizing the effectiveness (and even trust) of recommender systems. The task definition discussed above primarily expresses the inputs that a recommender model should consider and the desired output. However, the various costs incurred by users at different stages are not explicitly represented in the problem definition, but can be incorporated into the loss function design in the proposed models.
Table 1. Example recommendation tasks across different domains. The characteristics are based on our best understanding, which may not be applicable to any specific platform. The “price” attribute is ignored from pre-interaction judgment.
Domain Item Type Session Candidate Item Selection Pre-Interaction Judgment User-Item Interaction
Music streaming Single Yes Repeat, Exploration No, except the first song Progress ratio, Skip
Short video Single Yes No, or non-repeat No, except the first video Progress ratio, Watching speed, Skip
Online shopping Multi Might Repeat, Exploration Description, Image, Review Payment, Return
News Single Yes Region, Genre Title, Source Click, View
Job Single No Location Description Unobservable, Post-interaction inference
Course Single No Meeting pre-requisite Description, Instructor, Institute Progress, completion
Food delivery Single No Repeat, Exploration, Delivery distance/time Image, Description, Review Payment
POI Multi Might Region, Repeat, Exploration Description, Image, Travel cost, Review Unobservable, Post-interaction inference
5.2. Task, Solution, and Evaluation
A task formulation is often a formal abstraction of real-world applications. While such abstractions may omit details specific to certain recommendation platforms, the solutions developed should remain confined to the defined input and output of the task formulation. Consequently, since offline evaluation often serves as a proxy for selecting the most promising solutions for online evaluation, it should be designed to assess a solution’s performance with respect to the task formulation itself.
In RecSys research, evaluation is sometimes tailored to fit the proposed solutions rather than being aligned with the task formulation. One example is the evaluation of Bandits and reinforcement learning-based RecSys solutions. As reported in (Sun, 2023), when examining dataset partitioning methods, it was observed that only a small number of papers followed the timeline and simulated user interactions over time, which is a good practice. However, the evaluation was not motivated by the task itself, but because reinforcement learning solutions require reward signals based on user actions, leading to an evaluation setup that caters to the proposed solution.
Some existing RecSys tasks are not well formulated. Sequential recommendation and intent-aware recommendation are two examples; they do not fundamentally differ from a typical recommendation problem. In the survey paper, Jannach and Zanker (2024) defines: intent-aware recommender “is a recommender system that is designed to capture the users’ underlying current motivations and goals in order to support them.” If we view the problem definition by its inputs and outputs, there is no significant difference from the definition in Equation 6. Probably due to a less distinctive problem formulation, there are questions on the progress made (Shehzad et al., 2025). In the survey paper (Pan et al., 2024), the task of sequential recommendation is defined as follows: “In sequential recommendation, we often have one or more sequences of interacted items w.r.t. each user, as well as some auxiliary information to help learn user preferences. Our goal is then to generate a ranked list of items accurately for each user”. This is basically a different view of the very same $U\times I$. Naturally all items interacted with by a user form a sequence as illustrated in Figure 2. Solutions that pay attention on such sequences may be able to achieve a better recommendation accuracy, particularly in the RecSys settings with items of a single type e.g., music, movie, book, and news. However, the task to be addressed remains a generic recommendation.
A comparable research challenge in the past was imbalanced classification. In that domain, all classification problems with imbalanced class distributions were broadly categorized as “imbalanced classification problems,” and researchers sought a common solution to address the imbalance while often overlooking the true underlying reasons for it. However, data imbalance can arise from multiple causes, and while imbalance is an easily observable characteristic of data distribution, it is not itself the root cause (López et al., 2013). As a result, solutions designed specifically for “imbalanced classification”, the surface form of the data imbalance, may or may not be effective, leading to inconclusive results.
The proposed task-centric framework provides a set of dimensions to consider and compare RecSys tasks. Accordingly, the proposed solutions can be evaluated in a fair and meaningful way.
5.3. Task Specificity vs Model Generalizability
Recommender systems represent a fascinating intersection between theoretical research and practical deployment. On the one hand, they are directly linked to real-world applications, where performance improvements can lead to tangible business benefits and enhanced user experiences. On the other hand, academic research often strives to develop generic models that generalize well across diverse datasets and application scenarios. The tension between these two objectives creates an interesting challenge: we need models that perform well across multiple datasets representing different recommendation settings, yet optimizing a model for a specific application often requires customization in terms of input features, objective functions, and even model architecture.
This trade-off between task specificity and model generalizability is a fundamental issue in recommender system research. Highly specialized models, fine-tuned for a particular domain, can achieve state-of-the-art performance in their respective tasks but may struggle when applied to other recommendation settings. Conversely, more general models that are designed to work across various domains often sacrifice performance in any given task, as they cannot fully exploit the domain-specific characteristics that drive user preferences. This issue is further exacerbated by the diverse nature of recommendation tasks, as discussed earlier.
One potential approach to addressing this challenge is to establish a broad categorization of recommendation tasks based on key dimensions, such as candidate item selection, single- vs. multi-category recommendations, the cost and information available during pre-interaction judgment, the nature of user-item interactions, and the availability of post-interaction feedback. Table 1 provides an overview of several example recommendation domains and highlights their distinguishing characteristics.
From the table, we observe that music streaming and short-video recommendations share many common characteristics, except that users rarely prefer to repeatedly watch the same short videos. However, both domains differ significantly from POI (Point of Interest) recommendation. In POI recommendation, items include restaurants, parks, and historical landmarks, making it a multi-category recommendation task. Additionally, there is a fundamental distinction between recommending previously visited places (e.g., on a map or navigation interface) and suggesting new locations for exploration. This contrast underscores the importance of defining recommendation tasks based on their inherent characteristics rather than treating them as a single unified problem.
Such categorization ensures that all solutions are fairly evaluated on benchmark datasets regardless of the techniques used. We believe that, a detailed understanding and categorization of research tasks can provide insights into selecting appropriate baselines (Ivanova et al., 2023) and datasets (Beel et al., 2024). Ultimately, the effectiveness of a recommender system depends not necessarily on the sophistication of the model but on how well it aligns with the specific characteristics of the recommendation task at hand. Striking the right balance between generalizability and task specificity remains an open challenge and a key research direction for advancing the field.
By categorizing recommendation tasks along these dimensions, we provide a structured framework for understanding the diversity of recommender system applications. Future work should focus on refining these categorizations and developing standardized evaluation protocols that account for the unique characteristics of each recommendation task.
6. Conclusion
While foundational works in recommender systems have been well established, there remains a lack of consensus within the community regarding baseline models and datasets. This gap is likely due to an insufficient understanding of RecSys tasks. In this paper, we provide an in-depth analysis of recommender system task formulations, emphasizing the need for clear and well-defined problem definitions to support effective evaluation and the development of more applicable solutions. We stress the importance of understanding input-output relationships in recommendation models, considering factors such as temporal dynamics and candidate item selection. Additionally, we examine the complexities of user-item interactions, including user-incurred costs during decision-making and the challenges posed by multi-step interactions in real-world scenarios. We also emphasize the goal of RecSys is to minimum (all forms of) user costs which may occur at any stage.
Many points discussed here may have been mentioned in previous literature or are common practices in the industry. Our goal, as a perspective paper, is not to provide a comprehensive survey. Instead, we aim to present a cohesive framework for better characterizing RecSys tasks, thereby guiding future academic research toward more meaningful contributions. This also calls for more details to be included in datasets released by the industry, as offline evaluations are particularly important for academic researchers who do not have access to A/B testing.
Ultimately, this paper seeks to clarify the relationship between task formulation, solution development, and evaluation (more specifically offline evaluation). By fostering a more structured and comprehensive understanding of these key elements, we hope to provide valuable insights for both new and experienced researchers, contributing to the continued advancement of the field of recommender systems across diverse real-world domains.

### v3

@@
-Are We Solving a Well-Defined Problem? A Task-Centric Perspective on Recommendation Tasks
+A Task-Centric Perspective on Recommendation Systems
 Abstract.
-Recommender systems (RecSys) leverage user interaction history to predict and suggest relevant items, shaping user experiences across various domains.
-While many studies adopt a general problem definition, i.e., to recommend preferred items to users based on past interactions, such abstraction often lacks the domain-specific nuances necessary for practical deployment.
-However, models are frequently evaluated using datasets from online recommender platforms, which inherently reflect these specificities.
+Many studies in recommender systems (RecSys) adopt a general problem definition, i.e., to recommend preferred items to users based on past interactions.
+Such abstraction often lacks the domain-specific nuances necessary for practical deployment.
+However, models are frequently evaluated using datasets collected from online recommender platforms, which inherently reflect domain or task specificities.
 In this paper, we analyze RecSys task formulations, emphasizing key components such as input-output structures, temporal dynamics, and candidate item selection.
 All these factors directly impact offline evaluation.
-We further examine the complexities of user-item interactions, including decision-making costs, multi-step engagements, and unobservable interactions, which may influence model design and loss functions.
+We further examine the complexities of user-item interactions, including decision-making costs, multi-step engagements, and unobservable interactions, which may influence model design.
 Additionally, we explore the balance between task specificity and model generalizability, highlighting how well-defined task formulations serve as the foundation for robust evaluation and effective solution development.
@@
 Recommender Systems, Task Formulation, Evaluation, User cost
-copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Draft; March 2025; ACM Templateisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
+copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systems
 1. Introduction
@@
 The mismatch between abstract problem definitions and domain-specific evaluations leads to inconsistent settings and findings across experiments (Zangerle and Bauer, 2022; McElfresh et al., 2022).
-In this paper, we present a comprehensive analysis of different RecSys tasks, highlighting their key distinctions and implications for evaluation and real-world deployment.
-Recommender systems are not a new research area, and numerous influential works exist.
-We begin our discussion in Section 2 with a review of several highly cited works in RecSys, emphasizing task formulations.
+We begin with a review of several highly cited works in RecSys, emphasizing task formulations.
 Interestingly, key factors in RecSys were well defined and discussed decades ago.
 Yet, the community still disagrees on the choice of baselines and datasets (Ivanova et al., 2023; Beel et al., 2024).
-One main contribution of this paper is a detailed examination of task definition in Section 3, focusing primarily on the input and output of a mapping function, i.e., a recommender.
-We discuss a missing element in task formulation i.e., time, and the selection of candidate items for recommendation, both of which have a significant impact on offline evaluation.
-Another contribution of this paper is a detailed analysis of the life cycle of user-item interactions in Section 4.
-In the pre-interaction judgment stage, users are presented with a list of choices or recommended items.
-Users incur costs when making informed (or sometimes uninformed) decisions to select a recommendation, which serves as a clear indication of user preferences.
-We then explore the complexity of user-item interactions.
-For example, some interactions can involve multiple steps, as in e-commerce, where a user may add an item to their cart, proceed to payment, await delivery, interact with the product, and possibly return it.
-In music streaming, users are not offered a choice but can only skip or fastforward a recommended song.
-In some cases, true user-item interactions may not be directly observable as in job recommendation.
-Based on the task definition and the framework for understanding it, we present our perspectives in Section 5.
-We consider (i) the ultimate goal of recommendation, i.e., reducing user cost, (ii) the relationship between task formulation, solution, and evaluation, and (iii) the intersection of task specificity and model generalizability.
-In our discussion, we use a few examples to show the differences in RecSys tasks based on the presented dimensions, i.e., candidate item selection, pre-interaction judgment, and user-item interaction.
-We hope that this paper provides researchers, particularly those new to recommender systems, with a clearer understanding of the research topic.
-After establishing a clear understanding, subsequent research can precisely define specific RecSys tasks, select appropriate baselines, and design experiments that account for task specificity.
+We then provide a detailed examination of task definition, focusing primarily on the input and output of a mapping function, i.e., the recommender.
+We discuss the missing elements in task formulation: time and the selection of candidate items for recommendation, and their impact on offline evaluation.
+Lastly, we review the life cycle of user-item interactions, with a focus on the cost incurred from recommendations made to the feedback of the user-item interactions.
+Based on the task definition and proposed framework, we outline key perspectives and actionable directions for future work.
 2. A Historical Review of Task Formulation
-We begin with four of the most widely cited papers, [^1] Citation counts were obtained via a Google Scholar search for papers with the title words “recommender systems” or “recommender system” on February 12, 2025.
-each with over 8,000 citations.
+We begin with a few widely cited RecSys papers (Herlocker et al., 2004; Koren et al., 2009; Ricci et al., 2011; Adomavicius and Tuzhilin, 2005).
 As foundational works in this field, these papers have influenced many researchers, and the tasks they define have likely shaped numerous follow-up studies.
-The four papers reviewed were published between 2004 and 2011 (Herlocker et al., 2004; Koren et al., 2009; Ricci et al., 2011; Adomavicius and Tuzhilin, 2005).
-The Recommendation Problem. In one of the most influential survey papers in RecSys, Adomavicius and Tuzhilin (2005) formally define the recommendation problem as follows:
+In an influential survey paper, Adomavicius and Tuzhilin (2005) formally define the recommendation problem as follows:
 Definition 0 (Recommendation Problem).
@@
 The authors further note that “the utility of an item is usually represented by a rating, which indicates how a particular user liked a particular item,” a concept commonly known as explicit feedback.
-Nevertheless, in most subsequent RecSys studies, implicit feedback has become far more prevalent (Rendle et al., 2009).
-As a result, the utility of an item to a user is often inferred from binary feedback, whether the user has interacted with the item, such as listening to a song, purchasing a watch, or staying at a hotel.
+In subsequent RecSys studies, implicit feedback has become far more prevalent (Rendle et al., 2009).
+As a result, the utility of an item to a user is often inferred from binary feedback, whether the user has interacted with the item.
 Notably, the difference between explicit and implicit feedback primarily affects data modeling and the loss function design when learning $r$, while the core recommendation problem remains unchanged.
-Collaborative Filtering.
 Koren et al.
-(2009) do not formally define tasks in recommender systems but instead compare the two primary approaches: content filtering and collaborative filtering.
-They describe content filtering as a method that creates user or item profiles to capture their characteristics, enabling the recommender to associate users with relevant items.
-In contrast, collaborative filtering relies exclusively on past user-item interactions, such as previous transactions or product ratings.
-The authors highlight that “a major appeal of collaborative filtering is that it is domain-free, yet it can address data aspects that are often elusive and difficult to profile using content filtering,” further noting that collaborative filtering is “generally more accurate than content-based techniques.”
-Due to these advantages, collaborative filtering solutions are often considered more generalizable, aligning well with one of the key objectives in academic research.
-However, the authors also acknowledge the cold start problem in collaborative filtering, which arises when new users or items have no past interactions (Schein et al., 2002).
+(2009) compare the two primary approaches in RecSys: content filtering and collaborative filtering.
+Content filtering builds user and item profiles based on their characteristics, allowing the system to match users with relevant items.
+In contrast, collaborative filtering relies exclusively on past user-item interactions.
 Although solutions such as content-based, collaborative, and hybrid filtering are independent of problem definitions, the widespread adoption of collaborative filtering in recent research leads us to assume that user-item interactions remain a key input for typical recommender systems.
-With this in mind, we reiterate the points made in (Koren et al., 2009), emphasizing that collaborative filtering relies exclusively on past user-item interactions, such as previous transactions or product ratings.
-User Feedback and Preference.
-In the introduction chapter of the RecSys Handbook, Ricci et al.
-(2011) describe personalized recommendations as ranked lists of items, where recommender systems aim to “predict what the most suitable products or services are, based on the user’s preferences and constraints.” Regarding user preferences, the authors highlight that these can also be inferred from their actions, such as navigating to a specific product page.
+Regarding user preferences, Ricci et al.
+(2011) highlight that these can also be inferred from their actions, such as navigating to a specific product page.
 In this context, user feedback can take multiple forms: explicit feedback, such as ratings; implicit feedback, derived from user-item interactions; and inferred preferences based on observed user behavior.
-Task Diversity.
 Herlocker et al.
 (2004) provide a thorough discussion on user tasks for recommender systems.
-Their discussion focuses on end-user tasks (i.e., not marketers or other system stakeholders), [^2] We refer readers to (Zangerle and Bauer, 2022) for more detailed discussion on other users in a recommender system like item provider, platform provider, and other stakeholders.
-which aligns well with the RecSys tasks to be discussed in this paper.
+Their discussion focuses on end-user tasks (i.e., not marketers or other system stakeholders) [^1] We refer readers to (Zangerle and Bauer, 2022) for more detailed discussion on other users in a recommender system like item provider, platform provider, and other stakeholders., which aligns well with the RecSys tasks to be discussed in this paper.
 The key user task is to find good items, such as providing users with a ranked list of recommended items.
-The authors highlight that “there are likely to be many specializations of the tasks within each domain,” and the domain-specific characteristics are reflected in the properties of the datasets, e.g., the need for novelty versus quality, as well as the cost/benefit ratio of false/true positives/negatives.
-Regarding novelty, in certain domains, “the user goal is dominated by finding recommendations for things a user does not already know about.” In other domains, users expect recommendations to help them find high-quality items among many choices.
-Regarding cost/benefit, the authors provide an example of the relatively low cost of false-positive movie recommendations.
-However, in other settings, e.g., if an item comes with a high price tag, then false positives may carry a much higher cost.
-Summary.
-From this historical review, we derive the following main points: (i) The primary task of a recommender system is to suggest items to users based on their preferences.
-(ii) Collaborative filtering is a widely adopted approach that leverages past user-item interactions.
-(iii) In most cases, users provide implicit feedback, and their preferences can also be inferred from their actions.
-(iv) Recommendation tasks can be domain-specific, with various factors to consider, such as the desired level of novelty and quality, as well as the trade-off between cost and benefit.
+The authors highlight that “there are likely to be many specializations of the tasks within each domain,” and the domain-specific characteristics are reflected in the properties of the datasets.
 3. A Closer Look at the Task Definition
-Based on the above discussion and the main points identified, we now examine the RecSys task definition from multiple perspectives.
-For ease of presentation, we rewrite Definition 2.1 by specifying recommender as a mapping from its input to output.
-Figure 1.
-The classic task definition of recommendation, where user-item interactions are represented in a matrix form.
-The objective is to predict the missing values, which correspond to unobserved user-item interactions.
+For clarity, we rewrite Definition 2.1 by specifying a recommender as a mapping function.
 Definition 0 (Recommendation).
-Let $U$ be a set of users and $I$ be a set of items. A recommender system aims to produce a ranked list of items for a user $u$, based on user-item interactions $U\times I$.
+Let $U$ be a set of users and $I$ be a set of items. A recommender aims to produce a ranked list of items for a user $u$, based on user-item interactions $U\times I$.
 (2) $\langle u,U\times I,I\rangle\rightarrow R^{u}$
-In this rewritten definition, the recommender system’s input consists of three components: (i) the user $u$ for whom recommendations are to be made, (ii) the set of user-item interactions $U\times I$, which includes all existing interactions made by all users and all items, and (iii) the set of available items $I$ from which recommendations are being made.
+In this rewritten form, a recommender’s input consists of three components: (i) the user $u$ for whom recommendations are to be made, (ii) the set of user-item interactions $U\times I$, which includes existing interactions made by users and items, and (iii) the set of available items $I$ from which recommendations are being made.
 The system outputs a ranked list of recommended items for $u$, denoted by $R^{u}$.
-Following the tradition of illustrating recommender system tasks in matrix form, we present Definition 3.1 in Figure 1.
-In this figure, the goal is to generate a ranked list of recommended items for $u_{3}$.
-The items $i_{1}$ to $i_{n}$ collectively form the item set $I$, and the interactions encompass all interactions made between the $m$ users and $n$ items.
-Note that this definition is generic to cover all attributes or side information of users or items, because all such information or metadata can be easily derived given a user ID or item ID.
-As we focus on the core problem definition of RecSys in this paper, we omit all these details.
+This definition is generic to cover all attributes or side information of users or items, because all such information can be easily derived from user IDs and item IDs.
 3.1. The Missing Input: Time
-Recall that in (Koren et al., 2009), the authors note that collaborative filtering relies exclusively on past user-item interactions to make recommendations.
-In practical scenarios, whether recommending a product for purchase or a song to listen to, it is common sense that recommendations made at a time point $t$ should be based on all information available at or before $t$.
-Ideally, the user will find the recommendations appealing, then choose to interact with them after the new recommendations are made at $t$.
-Adding Time to the Input.
+In practical scenarios, whether recommending a product for purchase or a song to listen to, recommendations made at a time point $t$ should be based on all information available at $t$.
 With that, we rewrite the mapping function in Definition 3.1 to consider the time dimension.
@@
 (3) $\langle u,t,(U\times I)_{\leq t},I_{\leq t}\rangle\rightarrow R_{t}^{u}$
-In this rewritten form, we introduce a time point $t$, indicating a recommendation is to be made for user $u$ at time point $t$.
+Here, we introduce a time point $t$, indicating the time a recommendation is to be made for user $u$.
 We use $(U\times I)_{\leq t}$ to represent all user-item interactions that occurred before time $t$, a simplified form of $(u,i,t_{x})\in U\times I|t_{x}\leq t$}.
-Accordingly, the candidate items available for recommendation are those that were registered in the system and accessible to the recommender as at $t$, denoted by $I_{\leq t}$.
+The candidate items available for recommendation are those that were registered in the system and accessible as at $t$, denoted by $I_{\leq t}$.
 Figure 1.
-The recommendation task along time line.
 Recommendations are to be made for user $u_{2}$ at time point $t_{e}$.
-A model is expected to learn from interactions occurred before $t_{e}$ i.e., $(U\times I)_{\leq t_{e}}$ and make recommendation among items available at $t_{e}$, denoted by $I_{\leq t_{e}}$.
-While this reformulation may seem trivial and may not impact real-world or online RecSys recommendations, strictly adhering to this task definition poses significant challenges for offline evaluations (Sun, 2023).
-Impact to Offline Evaluation.
-In offline evaluations, user-item interactions in a dataset must be partitioned into training and test sets.
-The former is used to train the RecSys model, while the latter is used to assess its performance.
-According to the task formulation in Equation 3, evaluating a model’s performance at time $t$ for a user $u$ requires that the model only access or learn from the interactions that occurred before $t$, i.e., $(U\times I)_{\leq t}$, and generate recommendations from $I_{\leq t}$, the items available at or before $t$.
+A model is expected to learn from interactions occurred before $t_{e}$ i.e., $(U\times I)_{\leq t_{e}}$ and recommend items available at $t_{e}$, denoted by $I_{\leq t_{e}}$.:
+While this reformulation may seem trivial and does not impact real-world or online RecSys implementations, strictly adhering to this task definition poses significant challenges for offline evaluations (Sun, 2023).
+In offline evaluations, user-item interactions in a dataset are partitioned into training and test sets.
+The former is used to train a recommendation model, while the latter is used to assess its performance.
 Figure 1 provides an illustration where the items interacted with by each user are plotted to the right of the user along a global timeline.
@@
 The recommender can learn from all user-item interactions that occurred before $t_{e}$, i.e., $(U\times I)_{\leq t_{e}}$.
-The items available for recommendation to $u_{2}$ at $t_{e}$ are those that have been interacted with by any user before $t_{e}$.
+The items available for recommendation to $u_{2}$ at $t_{e}$ are those that have been interacted with by any user before $t_{e}$: $I_{\leq t_{e}}=\{t_{1},t_{2},t_{3},t_{4},t_{5}\}$.
 Notably, items $i_{7}$ and $i_{8}$ received their first interaction after $t_{e}$, the system has no interaction data for these items at $t_{e}$, and they should not be recommended to $u_{2}$.
 [^2] Note that our discussion assumes the recommendation model is based on collaborative filtering, to avoid the possible misunderstanding that items $i_{7}$ and $i_{8}$ could be recommended based on profile matching.
-Finally, note that the user-based leave-last-one-out data split results in each user having their own $(U\times I)_{\leq t}$ and $I_{\leq t}$, since not all users have their last interaction at a common time.
-Without enforcing a global time point, many widely adopted data partitioning schemes, such as random splitting or user-based leave-last-one-out, lead to data leakage (Ji et al., 2023).
-This occurs when user-item interactions from the future (relative to time $t$ in Equation 3 or $t_{e}$ in Figure 2) are used to train the model and make recommendations.
-In other words, when recommending items for a specific user, the model does not rely solely on that user’s interaction history but also leverages interactions from other users who may share similar preferences.
-The time point constrains all user-item interactions accessible to the system.
-The incorporation of factor $t$ into the task formulation also directly impacts the implementation of certain baseline models.
-For instance, the widely used popularity baseline, which simply counts item occurrences in the training data, does not accurately reflect real-life item popularity.
-In practice, item popularity is typically associated with a reference time point and a specific time range, e.g., the top-selling books from last week or last month.
-By introducing a time point $t$ in problem definition, item popularity can be derived with reference to that specific time point (Ji et al., 2020, 2022).
+Note that the user-based leave-last-one-out data split results in each user having their own $(U\times I)_{\leq t}$ and $I_{\leq t}$, since not all users have their last interaction at the same time.
+Without enforcing a global timeline or an absolute time point, many commonly used data partitioning schemes, like random splitting or user-based leave-last-one-out, can lead to data leakage.
+Empirically, we show that the impact of data leakage on recommendation models is unpredictable (Ji et al., 2023).
+Hence, comparing the performance of recommendation models under data leakage in offline evaluation is both impractical and meaningless.
+Recently, Le et al.
+(2025) further show that, for sequential recommenders, eliminating data leakage leads to a 21.7 - 73.4% drop in sampled nDCG@10 compared to the commonly adopted setting that includes data leakage.
+They also observe that the impact of data leakage on model rankings is unpredictable, consistent with our findings in (Ji et al., 2023).
+Indeed, some recent studies have realized the importance of this issue, and adopted splits based on absolute temporal cutoffs.
+To examine how prevalent this practice is, we reviewed the 49 long papers published at the ACM RecSys 2025 conference (September 2025).
+We found that 16 papers adopted random splits, entirely ignoring the temporal factor.
+Among the remaining papers that report experimental settings, 7 used relative temporal splits, which also lead to data leakage.
+The incorporation of the temporal factor $t$ into the task formulation also directly affects the implementation of certain baseline models.
+For example, the widely used popularity baseline (which simply counts item occurrences in the training data) does not accurately reflect real-world item popularity.
+In practice, item popularity is typically defined with respect to a reference time point and within a specific time window, e.g., the top-selling books from the past week or month.
+By explicitly introducing a time point $t$ into the problem definition, item popularity can be computed relative to that specific time point.
+Interestingly, the simple temporally aware popularity model, DecayPop (Ji et al., 2020), has been shown to be the most effective ranking method on the Yambda-5B dataset for the music recommendation ‘like’ scenario, outperforming many more complex models (Ploshkin et al., 2025).
+This finding calls for a careful re-evaluation of prior studies that overlooked the temporal dimension.
 3.2. The Missing Constraints on Candidate Items
-Next, we focus on the available items that can be recommended to a user $u$ at time $t$.
 Herlocker et al.
-(2004) discuss recommender system tasks from the perspective of novelty and quality, assuming that users generally expect recommended items to be new and previously unconsumed.
-However, this assumption may not always hold, as a recommender system is ultimately designed to help users find preferred items more easily and quickly.
-In many scenarios, users may prefer to interact with items they are already familiar with and confident in (Anderson et al., 2014).
+(2004) discuss recommendation tasks from the perspective of novelty and quality, assuming that users generally expect recommended items to be new and previously unconsumed.
+In some scenarios, users may prefer to interact with items they are already familiar with and confident in (Anderson et al., 2014).
 Groceries shopping (Ariannezhad et al., 2022; Katz et al., 2022), e-commerce (Wang et al., 2019; Quan et al., 2023), food delivery ordering (Li et al., 2024), and song listening (Tsukuda and Goto, 2020; Reiter-Haas et al., 2021) are a few examples where a user may prefer to have earlier consumed items to be recommended for easy selection.
 Let $I^{u}_{t}$ be the items that $u$ has interacted with in the past at time point $t$.
-We use $I_{t}^{\bar{u}}$ to denote the remaining items available at $t$ that are new to $u$, then the recommendation in Equation 3 can be divided into two formulations for repeated consumption $R_{t}^{ur}$ and for exploration $R_{t}^{ue}$ recommendations, respectively.
+We use $I_{t}^{\bar{u}}$ to denote the remaining items available at $t$ that are new to $u$.
+Equation 3 can be divided into two formulations for repeated consumption $R_{t}^{ur}$ and for exploration $R_{t}^{ue}$ recommendations, respectively.
 (4) $\langle u,t,(U\times I)_{\leq t},I_{\leq t}^{u}$ $\rangle\rightarrow R_{t}^{ur}$
 (5) $\langle u,t,(U\times I)_{\leq t},I_{\leq t}^{\bar{u}}$ $\rangle\rightarrow R_{t}^{ue}$
-Note that, how to present $R_{t}^{ur}$ and $R_{t}^{ue}$ to users through a web or app interface, e.g., as two separate rankings or as a merged ranking of items from both recommendations, is orthogonal to our discussion in this paper.
-Figure 3.
-The recommendation for repeated consumption for user $u_{2}$ at time $t_{e}$.
-The search space for repeated consumption is the two items $i_{2}$ and $i_{5}$.
-To recommend new items, the candidate items will be the remaining items in $I_{\leq t_{e}}\setminus\{i_{2},i_{5}\}$.
-Impact to Offline Evaluation.
-Again, the partitioning of candidate items for recommendation $I$ into $I^{u}$ and $I^{\bar{u}}$ may seem trivial (here, we omit the subscript $t$ for clarity), but this partitioning leads to (i) significant differences in the search space for a recommender, because $|I^{u}|\ll|I^{\bar{u}}|$ and $|I^{\bar{u}}|\approx|I|$, and (ii) subsequently significant differences or impacts on evaluation.
-It is important to note that if repeated consumption and exploration are not clearly distinguished, the evaluation results may be significantly biased toward repeated consumption.
-Because of the much smaller item search space for recommendation and a large portion of, if not all, items being users’ preferred items, the chances of making a good recommendation among the repeated items are much higher than picking an item from a much larger pool of items, especially those the user has no experience with.
-The recommendation accuracy could be easily dominated by the repeated items if repeated consumption is common like food ordering and groceries shopping.
-Hence, the separation of evaluations of repeated consumption and exploration better reflects the model’s accuracy; a detailed discussion and comparison is presented in  (Li et al., 2024) for food delivery recommendation.
-Additionally, due to the much smaller search space, recommending repeated items is significantly more efficient than recommending items for exploration.
-Candidate Item Selection.
+Note that, how to present $R_{t}^{ur}$ and $R_{t}^{ue}$ to users, e.g., as two separate rankings or as a merged ranking of items from both recommendations, is orthogonal to our discussion here.
+The partitioning of candidate items into $I^{u}$ and $I^{\bar{u}}$ may again seem trivial (the subscript $t$ is omitted for clarity), but it leads to (i) significant differences in the recommender’s search space, since $|I^{u}|\ll|I^{\bar{u}}|$ and $|I^{\bar{u}}|\approx|I|$, and (ii) consequently, notable impacts on evaluation.
+Because the recommendation space is much smaller and largely composed of earlier preferred items, making good recommendations among repeated items is far easier than selecting from a large pool of unfamiliar ones.
+For instance, in food delivery recommendation, the proposed model in (Li et al., 2024) achieved NDCG scores ranging from 0.59 to 0.64 for repeat consumption, while for exploration they ranged only from 0.09 to 0.17 across datasets from three cities.
+Similar performance gaps were observed for all evaluated baselines (Li et al., 2024).
+The recommendation accuracy could be easily dominated by the repeated items if repeated consumption is common in the task setting, like food ordering and groceries shopping.
+Hence, the separation of evaluations of repeated consumption and exploration better reflects the model’s accuracy.
 The repeated consumption and exploration is just one example of candidate item selection.
-The candidate items suitable for recommendation can be directly specified by users.
-For example, if a user is planning a trip to a city, hotel recommendations would only be meaningful if they are located in that city and available during the user’s travel dates.
-Although this may seem trivial and makes the recommendation problem more similar to a search problem, the vast majority of recommendation work using the Yelp dataset does not consider users’ geographical constraints, to the best of our knowledge.
-That is, if a user resides in New York City and primarily visits restaurants within a relatively small region of NYC, recommending a restaurant from a different state would be far less practical for this user.
-On the other hand, the candidate items suitable for recommendation could also be derived based on user intent mining through search logs and other means.
-Note that the user intent mining here is different from the line of research on intent-aware recommender systems (Jannach and Zanker, 2024; Shehzad et al., 2025), which will be discussed in Section 5.2.
-One key difference is that, in our discussion on candidate item selection, the set of eligible items to be recommended is determined before running the recommender system, hence is part of the input to the problem definition and can be specified by users, illustrated in Figure 4.
-3.3. Yet Another Rewritten Task Formulation
-We can now combine Equations 4 and 5 into a more general formulation by introducing a selection condition on the candidate items, denoted as $s(I_{\leq t})$.
-Specifically, $s(\cdot)$ can represent selection criteria based on user interaction history e.g., for repeated consumption, or be specified by users with temporal, spatial, or other attribute-based conditions.
-However, by listing $s(I_{\leq t})$ as part of the input to the recommendation task, the candidate set of items is determined before a model ranks them.
+The candidate items suitable for recommendation can be directly specified by users e.g., by time, location, or any other item attributes.
+The key message here is that the set of eligible items to be recommended is determined before running the recommender system.
+3.3. Task Formulation with Constraints
+We can now define a more general formulation by introducing a selection condition on the candidate items, denoted as $s(I_{\leq t})$.
+Specifically, $s(\cdot)$ represents selection criteria derived from a user’s interaction history like repeated consumption, or be specified by the user through temporal, spatial, or other attribute-based constraints.
+By including $s(I_{\leq t})$ as part of the task input, the candidate set of items is determined prior to model ranking.
+In fact, candidate generation (also known as recall) constitutes the first stage of the multi-stage recommender system architecture widely adopted by today’s largest online platforms, preceding the ranking and re-ranking stages (Liu et al., 2022).
 Definition 0 (Recommendation Task).
@@
 (6) $\left\langle u,t,(U\times I)_{\leq t},s(I_{\leq t})\right\rangle\rightarrow R_{t}^{u}$
-Now, we define the recommendation task as making a recommendation at time $t$ for user $u$ by referencing user-item interactions that occurred before $t$ and selecting suitable candidate items that satisfy the selection criteria $s(\cdot)$ to meet the inferred user preference.
-Following the example of repeat consumption and exploration, the final recommendation items presented to a user could be the results of running multiple recommendations made from multiple sets of candidate items conditioned on different multiple $s(\cdot)^{\prime}s$.
-Nevertheless, in terms of model evaluation, it is more meaningful to independently evaluate each model specific to candidate items on one selection function $s(\cdot)$.
+The final items recommended to a user could be the results of running multiple recommendations made from multiple sets of candidate items conditioned on different multiple $s(\cdot)$’s.
+In terms of model evaluation, it is more meaningful to independently evaluate each model specific to candidate items on one selection function $s(\cdot)$.
 4. From Recommendation to User Consumption
-Figure 4.
-User costs after recommendations made at time $t$: from pre-interaction judgment, to interaction, then to an optional post-interaction feedback.
-Among the recommended items, $i_{2}$ satisfies selection criteria $S(\cdot)$.
-So far, we have discussed the importance of explicitly incorporating the time factor into the problem definition and examined suitable candidate items for recommendation based on the context of the recommendation task.
-We consider these to be essential factors in the task definition due to their direct impact on model evaluation, which in turn directly impact the model design.
-Next, we take a closer look at user-item interactions $U\times I$ and analyze their meaning, as instances $(u,i,t_{x})$ are widely used to train recommender systems in academic research.
-There are many reasons for a user to interact with an item, such as through an intentional search rather than a recommendation.
-Since we are discussing recommendations, we make a general assumption that if a user interacts with an item, it reflects some form of preference for that item.
-However, we aim to explore the various stages from a recommended item $i\in R_{t}^{u}$ to an interaction $(u,i,t_{x})$, and to what extent a model can learn from the interaction.
+Users may interact with items for various reasons.
+An interaction could be the result of an intentional search rather than recommendation.
+For instance, the Yambda-5B dataset includes an is_organic flag to indicate actions not driven by recommendations (Ploshkin et al., 2025).
+In this discussion, we generally assume that a user’s interaction with an item reflects some degree of preference for it.
+Our focus is to examine the stages from a recommended item $i\in R_{t}^{u}$ to an interaction $(u,i,t_{x})$, and to what extent a model can learn from such interactions.
 This process can be considered as another form a mapping:
 (7) $\langle u,R_{t}^{u}\rangle\rightarrow(u,i,t_{x})\text{\quad where \quad}i\in R_{t}^{u}$
-4.1. User-Item Interaction: The Life Cycle
-In the early stages of RecSys research, many studies focused on explicit user feedback, such as ratings e.g., a user giving a movie a 5 out of 5 rating (Herlocker et al., 2004).
-However, since explicit feedback is often difficult to obtain, implicit feedback has become the mainstream in academic research e.g., a user has watched a movie.
-Nevertheless, Ricci et al.
-(2011) suggest that user preference can also be inferred from actions such as navigating to a product page e.g., a user clicking on a movie for more details among a list of candidate movies.
-However, there is a lack of detailed discussion on to what extent such feedback contribute to accurately modeling user preference.
-In our discussion, we divide user-item interactions into three stages: pre-interaction judgment, interaction, and post-interaction feedback, illustrated in Figure 4.
-While these stages may not apply to all recommendation scenarios, they provide a useful framework for understanding the relationship between user-item interactions and user preferences.
-Consider a user searching for a hotel in a city she has never visited, for a conference.
-She is presented with a list of 10 hotels on a booking website, initiating the pre-interaction judgment stage.
-She clicks on one hotel for more details, a decision influenced by factors such as images, branding, location, price, or other readily available attributes.
-Next, she enters the deliberate evaluation phase, where she reads the hotel description, checks room facilities, and reviews user feedback before deciding to book.
-All these effort happens before the user stays at the hotel.
-A few months later, she attends the conference and stays at the hotel.
-This marks the interaction stage, where she gains firsthand experience.
-Finally, after checking out, she provides feedback in the form of a rating and review, entering the post-interaction feedback stage.
-This feedback may consider various factors in a typical hotel review, such as location convenience, price, food, service quality, and hygiene.
+4.1. The Life Cycle of User-Item Interaction
+We divide user-item interactions into three stages: pre-interaction judgment, interaction, and post-interaction feedback.
+These stages may not apply to all recommendation scenarios, but offer a useful framework for understanding the relationship between user interactions and preferences.
+A conceptually similar framework was proposed by Lee and Kim (2023), who distinguish between pre-use preference (a user’s impression of an item before interacting with it) and post-use preference (the impression formed after interaction).
+Our discussion aims to highlight the different types of pre-interaction and interaction across various recommendation scenarios, which differs from both the technical solutions presented in (Lee and Kim, 2023) and the decision-making perspectives discussed in (Jameson et al., 2022).
+Consider a user booking a hotel in an unfamiliar city for a conference.
+When she first browses a list of options, she forms a pre-interaction judgment based on visible cues such as images, branding, location, and price.
+After clicking on one hotel, she enters the deliberate evaluation stage — reading descriptions, checking facilities, and reviewing feedback before booking.
+Months later, her stay in the hotel marks the interaction stage.
+After checkout, she leaves a rating and review, marking the post-interaction feedback stage.
 Both quick judgments and deliberate evaluations in the pre-interaction stage come at a cost, the user’s time and effort in gathering relevant information.
-The outcomes of these efforts strongly indicate user preference, especially in similar future situations, e.g., planning another trip or booking a restaurant.
+The outcomes of these efforts typically indicate user preference, especially in similar future situations, e.g., planning another trip.
 Post-interaction feedback, however, may not always accurately reflect user preference.
 While a review can highlight the primary reasons for choosing a hotel, representing genuine user preference, it may also describe aspects such as expectation gaps and booking effort, which are not necessarily preference indicators.
-A rating, influenced by these factors, may not always be a reliable measure of preference compared to the concrete fact that the user chose to stay at the hotel.
-In fact, in many practical scenarios, post-interaction feedback is often difficult to collect, as it requires users to put in additional effort to provide a rating or write a review.
+A rating, influenced by these factors, may not always be a reliable measure of preference compared to the fact that the user chose to stay at the hotel.
+In many practical scenarios, post-interaction feedback is often difficult to collect, as it requires users to put in additional effort.
 Hence, in our following discussion, we focus on the other two stages.
-In general, recommendations are presented as a list of candidate items before any user-item interaction occurs.
-At this stage, the user relies solely on their judgment, which inherently reflects preference, as they would likely make the same choice in similar circumstances.
-However, most academic datasets only contain actual interactions, lacking context on the reasons behind them.
-This limitation makes it challenging to model user preference purely based on observed interactions.
-Thus, a deeper understanding of the user decision-making process is crucial for more effective preference modeling.
 4.2. Complexity of Pre-Interaction Judgment
-We now focus on the pre-interaction judgment, which often takes the form of quick judgment or deliberate evaluation before a choice is made.
-The complexity of this judgment from the user’s perspective may arise from several dimensions.
-Informed vs Uninformed Decision.
+The complexity of pre-interaction judgment from the user’s perspective may arise from several dimensions.
+4.2.1. Informed vs Uninformed Decision
 One major factor from the user’s perspective is whether they possess the knowledge to accurately judge an item before interacting with it.
@@
 However, if a user has never used a robot vacuum before, many of the terms in the product description may be unfamiliar to them.
-Even after reading the machine specifications and user reviews, they may struggle to discern the pros and cons of a specific model, with references to her own home layout and floor conditions.
-These uninformed decisions may not necessarily indicate user preferences.
+Even after reading the machine specifications and user reviews, she may struggle to discern the pros and cons of a specific model, with references to her own home layout and floor conditions.
+Uninformed decisions may not necessarily indicate user preferences.
 It is also challenging for a recommendation platform to determine whether a user’s decision was based on prior knowledge or made without a full understanding of the item.
 This distinction is crucial, as recommendations based on uninformed interactions might not truly reflect user interests, potentially leading to less effective personalization.
-For example, after using a robot vacuum for some time, the user may realize that he should purchase another model with features better suited to his floor conditions.
-Items in One vs.
-Multiple Types/Categories.
+For example, after using a robot vacuum for some time, the user may realize that she should purchase another model with features better suited to her floor conditions.
+4.2.2. Items in One vs. Multiple Types
 Many studies in recommendation are conducted on datasets containing only one type of item, such as books, movies, news, or music.
 In these cases, there are common characteristics, such as genre, director, or artist, that users can rely on for pre-interaction judgment before actually interacting with a recommended item.
-However, in the context of online shopping, where products span thousands of categories, users apply different criteria and expectations when making judgments for different types of items, for informed decisions.
+However, in the context of online shopping, where products span thousands of categories, users apply different criteria and expectations when making judgments for different types of items.
 In such a diverse setting, user preferences learned from interactions across multiple item types may be shaped more by general associative patterns than by genuine preferences for specific products.
@@
 4.3. Recognition of User-Item Interaction
-After the pre-interaction judgment phase, the user makes a decision and interacts with an item.
-In most datasets used for RecSys evaluation, user-item interactions are recorded in the form of $(u,i,t_{x})$.
-However, interactions may appear in different forms depending on the recommendation context.
-Interaction Process Can Be Complicated.
+User-item interactions are often recorded in the form of $(u,i,t_{x})$ in a RecSys dataset. However, interactions may appear in different forms depending on the recommendation context.
+4.3.1. Interaction Process Can Be Complicated.
 Taking online shopping as an example, the process does not end when a user clicks on a recommended item.
@@
 This raises an important question: should an unsuccessful purchase (i.e., one that results in a return) still be considered a valid user-item interaction for learning user preferences?
-On one hand, the initial decision to purchase the item indicates some level of interest or preference.
+On the one hand, the initial decision to purchase the item indicates some level of interest or preference.
 On the other hand, the return suggests dissatisfaction or misalignment with the user’s expectations.
@@
 Therefore, when incorporating interaction data into preference modeling, it is crucial to differentiate between successful and unsuccessful purchases and consider additional contextual signals, such as return rates, to better understand user preferences.
-Absence of Pre-Interaction Judgment.
-Some of the user-item interactions happen without a pre-interaction judgment phase.
+4.3.2. Interaction without Pre-Interaction Judgment.
+Some user-item interactions happen without a pre-interaction judgment phase.
 This is especially common in scenarios like music streaming and short-video viewing, where users often do not actively select each item.
@@
 Unlike traditional recommendation settings where users consciously evaluate and select items before interaction, here, user preferences are inferred more dynamically based on real-time behaviors, or the interaction itself.
-In fact, users are less likely to explicitly express “like” in a streaming setting but are more likely to express “dislike”.
-This distinction raises important challenges for recommendation models, as they must distinguish between passive exposure and active preference while adapting to continuously evolving user interests.
+Recommendation models must distinguish between passive exposure and active preference while adapting to continuously evolving user interests.
 In this case, the common form of user-item interaction $(u,i,t_{x})$ becomes less accurate compared to other settings.
 Such recommendation in streaming form also post questions in item attribute modeling e.g., evaluating whether the cover image of a short video influence user viewing as the user may not even has the chance of viewing the cover image for each video in the streaming.
-Unobservable Interaction. There are also recommendation scenarios where user-item interactions are not directly observable and can only be inferred from external sources.
+4.3.3. Unobservable Interaction.
+There are also recommendation scenarios where user-item interactions are not directly observable and can only be inferred from external sources.
 One example is job recommendation, where matching is based on a user’s skills and knowledge against job requirements.
@@
 In the latter case, subsequent recommendations may be influenced by the previous selections or even the initial choice made at the start of the session.
-Figure 5. Reasons for a bad recommendation. The option “unrelated to the current video” suggests the dependency between the next recommendation and the current item.
-Figure 5 shows a feedback form for assessing video recommendation quality from YouTube.
-Notably, one of the options for a poor recommendation is labeled as “unrelated to the current video.” This indicates that the recommendation system tailors the next suggested video based on the one currently being played.
-The inclusion of this feedback option emphasizes users’ expectations that recommendations within a session should maintain contextual relevance.
 This is a key characteristic of session-based recommendation, a specialized type of recommendation task (Li et al., 2025).
 A detailed discussion on the recommendation flow is made in (Sun, 2024).
-This example further highlights the complexity of recommendation systems and the variety of factors that need to be accounted for when designing different recommendation tasks.
 Each recommendation scenario, whether based on user preferences, session context, or product types, requires careful formulation to ensure that the user experience is optimized and that the system accurately reflects user intent.
 5. Discussion and Perspectives
-Next, we discuss the ultimate goal of recommendation, i.e., reducing user cost, the relationship between task formulation, solution, and evaluation, and the intersection of task specificity and model generalizability, along with our perspectives on RecSys.
+Next, we discuss the ultimate goal of recommendation, examine the relationships among task formulation, solution, and evaluation, explore the intersection of task specificity and model generalizability, and provide actionable guidance for RecSys academic research.
 5.1. Recommendation vs User Cost
@@
 However, users still incur different costs at various stages of the interaction process.
-In the pre-interaction judgment stage, users must evaluate recommendations based on available information, such as descriptions, images, reviews, or other metadata, which requires cognitive effort and time.
-This cost is higher when users lack prior knowledge of the item category or when the recommendation list is overwhelming.
+In the pre-interaction judgment stage, users evaluate recommendations based on available information, such as descriptions, images, reviews, or other metadata, which requires cognitive effort and time.
 During the interaction stage, users experience the actual product or service, which may involve monetary costs (e.g., purchasing a product), time investment (e.g., watching a recommended movie), or engagement effort (e.g., exploring an unfamiliar interface).
-If the recommendation is poor, users may feel frustrated, leading to dissatisfaction and disengagement.
-Notably, such negative experiences can accumulate over time, eventually causing users to leave the platform, confirmed by a recent large scale survey (Zou et al., 2024).
+If the recommendation is poor, users may feel frustrated, leading to dissatisfaction and disengagement (Zou et al., 2024).
 Finally, in the post-interaction stage, users may be asked to provide feedback, such as ratings or reviews.
-While valuable for possibly improving recommendation quality, this step requires additional effort, and many users may choose not to participate.
-Understanding and minimizing all such costs at each of the three stages is essential for enhancing user experience and optimizing the effectiveness (and even trust) of recommender systems.
-The task definition discussed above primarily expresses the inputs that a recommender model should consider and the desired output.
-However, the various costs incurred by users at different stages are not explicitly represented in the problem definition, but can be incorporated into the loss function design in the proposed models.
-Table 1.
-Example recommendation tasks across different domains.
-The characteristics are based on our best understanding, which may not be applicable to any specific platform.
-The “price” attribute is ignored from pre-interaction judgment.
-Domain Item Type Session Candidate Item Selection Pre-Interaction Judgment User-Item Interaction
-Music streaming Single Yes Repeat, Exploration No, except the first song Progress ratio, Skip
-Short video Single Yes No, or non-repeat No, except the first video Progress ratio, Watching speed, Skip
-Online shopping Multi Might Repeat, Exploration Description, Image, Review Payment, Return
-News Single Yes Region, Genre Title, Source Click, View
-Job Single No Location Description Unobservable, Post-interaction inference
-Course Single No Meeting pre-requisite Description, Instructor, Institute Progress, completion
-Food delivery Single No Repeat, Exploration, Delivery distance/time Image, Description, Review Payment
-POI Multi Might Region, Repeat, Exploration Description, Image, Travel cost, Review Unobservable, Post-interaction inference
+This step requires additional effort, and many users may choose not to participate.
+Understanding and minimizing these costs at each of the three stages is essential for improving user experience and optimizing the effectiveness, and even the trustworthiness, of recommender systems.
+However, as illustrated in the examples above, such costs manifest differently across recommendation scenarios.
+The RecSys task definition in Equation 6 primarily specifies the inputs a recommender model should consider and the desired output.
+Yet, the various costs incurred by users at different stages from recommendation to interaction (Equation 7) cannot be explicitly represented in a formal formulation.
+Nevertheless, understanding these costs can inform the design of more refined loss functions for learning recommendation models.
+For example, in short-video recommendation, videos that a user watches for a duration shorter than their average viewing percentage can be treated as tolerance samples.
+In (Zou et al., 2024), such samples are assigned different losses compared to fully watched videos, leading to improvements in user retention verified through A/B testing.
+In this context, watching an uninteresting video constitutes a cost to the user.
 5.2. Task, Solution, and Evaluation
@@
 Probably due to a less distinctive problem formulation, there are questions on the progress made (Shehzad et al., 2025).
-In the survey paper (Pan et al., 2024), the task of sequential recommendation is defined as follows: “In sequential recommendation, we often have one or more sequences of interacted items w.r.t.
+In the survey paper (Pan et al., 2026), the task of sequential recommendation is defined as follows: “In sequential recommendation, we often have one or more sequences of interacted items w.r.t.
 each user, as well as some auxiliary information to help learn user preferences.
@@
 However, the task to be addressed remains a generic recommendation.
-A comparable research challenge in the past was imbalanced classification.
-In that domain, all classification problems with imbalanced class distributions were broadly categorized as “imbalanced classification problems,” and researchers sought a common solution to address the imbalance while often overlooking the true underlying reasons for it.
-However, data imbalance can arise from multiple causes, and while imbalance is an easily observable characteristic of data distribution, it is not itself the root cause (López et al., 2013).
-As a result, solutions designed specifically for “imbalanced classification”, the surface form of the data imbalance, may or may not be effective, leading to inconclusive results.
-The proposed task-centric framework provides a set of dimensions to consider and compare RecSys tasks. Accordingly, the proposed solutions can be evaluated in a fair and meaningful way.
 5.3. Task Specificity vs Model Generalizability
@@
 This issue is further exacerbated by the diverse nature of recommendation tasks, as discussed earlier.
-One potential approach to addressing this challenge is to establish a broad categorization of recommendation tasks based on key dimensions, such as candidate item selection, single- vs.
-multi-category recommendations, the cost and information available during pre-interaction judgment, the nature of user-item interactions, and the availability of post-interaction feedback.
-Table 1 provides an overview of several example recommendation domains and highlights their distinguishing characteristics.
-From the table, we observe that music streaming and short-video recommendations share many common characteristics, except that users rarely prefer to repeatedly watch the same short videos.
-However, both domains differ significantly from POI (Point of Interest) recommendation.
-In POI recommendation, items include restaurants, parks, and historical landmarks, making it a multi-category recommendation task.
-Additionally, there is a fundamental distinction between recommending previously visited places (e.g., on a map or navigation interface) and suggesting new locations for exploration.
-This contrast underscores the importance of defining recommendation tasks based on their inherent characteristics rather than treating them as a single unified problem.
-Such categorization ensures that all solutions are fairly evaluated on benchmark datasets regardless of the techniques used.
-We believe that, a detailed understanding and categorization of research tasks can provide insights into selecting appropriate baselines (Ivanova et al., 2023) and datasets (Beel et al., 2024).
-Ultimately, the effectiveness of a recommender system depends not necessarily on the sophistication of the model but on how well it aligns with the specific characteristics of the recommendation task at hand.
-Striking the right balance between generalizability and task specificity remains an open challenge and a key research direction for advancing the field.
-By categorizing recommendation tasks along these dimensions, we provide a structured framework for understanding the diversity of recommender system applications.
-Future work should focus on refining these categorizations and developing standardized evaluation protocols that account for the unique characteristics of each recommendation task.
+5.4. From Theory to Practice
+Moving forward, a practical step toward addressing the challenges discussed earlier is to establish a clear categorization of recommendation tasks.
+In our recent survey (Zou and Sun, 2025), which includes only research papers reporting online A/B testing results from production environments, we classify real-world recommendation tasks into two main categories.
+Transaction-oriented recommender systems generate item recommendations with the primary goal of prompting transactional user actions, optimizing for metrics such as conversion rate, revenue, or purchase likelihood.
+E-commerce platforms are typical examples.
+Content-oriented recommender systems, on the other hand, generate recommendations to facilitate user consumption and engagement, optimizing for metrics such as dwell time, clicks, or user satisfaction.
+Common examples include news, video, and music recommenders.
+Although these two main categories do not account for all types of recommendation systems, they capture the majority of widely studied recommendation scenarios.
+The objectives of recommender systems across these two broad categories differ substantially and can vary further within each category.
+Consequently, more specific recommendation tasks can be refined along key dimensions discussed earlier.
+For instance, recency is a key factor in news recommendation but may be less relevant for other types of content-oriented recommendation.
+RecSys research focused on algorithmic advancement should clearly specify the recommendation task it addresses, use datasets and evaluation settings that reflect realistic conditions, and compare against baselines designed for similar contexts.
+Such categorization ensures fair and meaningful evaluation across different recommendation tasks.
+Importantly, this emphasis on task specificity also requires a shared understanding within the community, especially among reviewers, that researchers should not be expected to overgeneralize their solutions to unrelated datasets or problem settings.
+For example, a submission should not be penalized for not reporting results on widely used datasets like MovieLens, when it addresses fundamentally different recommendation scenarios.
+The effectiveness of a recommender system depends much on how well it aligns with the specific characteristics of the target task.
+We highlight two recent examples that suggest researchers and reviewers are increasingly moving in this direction.
+Charolois-Pasqua et al.
+(2025) propose a playlist-generation recommender focused on a specific task: generating playlists from titles.
+Their method is evaluated on the Million Playlist Dataset and compared against task-relevant baselines.
+The authors further analyze user effort in playlist creation and assess the usefulness of playlist titles, demonstrating strong task alignment rather than pursuing cross-domain generalization.
+Similarly, based on their experience building a small-scale production news recommender system, Higley et al.
+(2025) observe that “published models are surprisingly difficult to apply to the kinds of data found in real-world datasets and practical recommendation problems.” They emphasize that “building recommender systems that serve real users requires deep and specific engagement not just with a domain in general, but with the particular characteristics of specific datasets, applications, and user communities,” underscoring the task-specific nature of RecSys research.
+Task-specific RecSys research has become increasingly feasible thanks to dataset availability.
+For example, Yambda-5B includes both implicit (e.g., listening) and explicit (e.g., likes and dislikes) feedback, as well as organic user actions specific to the music streaming recommendation task (Ploshkin et al., 2025).
+The authors also introduce a global temporal split to better reflect real-world settings and prevent data leakage.
 6. Conclusion
-While foundational works in recommender systems have been well established, there remains a lack of consensus within the community regarding baseline models and datasets.
-This gap is likely due to an insufficient understanding of RecSys tasks.
-In this paper, we provide an in-depth analysis of recommender system task formulations, emphasizing the need for clear and well-defined problem definitions to support effective evaluation and the development of more applicable solutions.
-We stress the importance of understanding input-output relationships in recommendation models, considering factors such as temporal dynamics and candidate item selection.
-Additionally, we examine the complexities of user-item interactions, including user-incurred costs during decision-making and the challenges posed by multi-step interactions in real-world scenarios.
-We also emphasize the goal of RecSys is to minimum (all forms of) user costs which may occur at any stage.
-Many points discussed here may have been mentioned in previous literature or are common practices in the industry.
-Our goal, as a perspective paper, is not to provide a comprehensive survey.
-Instead, we aim to present a cohesive framework for better characterizing RecSys tasks, thereby guiding future academic research toward more meaningful contributions.
-This also calls for more details to be included in datasets released by the industry, as offline evaluations are particularly important for academic researchers who do not have access to A/B testing.
-Ultimately, this paper seeks to clarify the relationship between task formulation, solution development, and evaluation (more specifically offline evaluation).
-By fostering a more structured and comprehensive understanding of these key elements, we hope to provide valuable insights for both new and experienced researchers, contributing to the continued advancement of the field of recommender systems across diverse real-world domains.
+While foundational works in recommender systems are well established, there remains a lack of consensus within the community regarding baseline models and datasets.
+The reason behind is likely stemming from an insufficient understanding of RecSys task formulation.
+In this paper, we provide an in-depth analysis of recommendation tasks, emphasizing the need for clear and well-defined problem definitions to enable effective evaluation and the development of more practical solutions.
+We highlight the importance of understanding input-output relationships in recommendation models, accounting for factors such as temporal dynamics and candidate item selection.
+We further examine the complexities of user-item interactions, including user-incurred costs during decision-making and the challenges introduced by multi-step interactions in real-world settings.
+Ultimately, we argue that the central objective of RecSys is to minimize user costs across the entire recommendation process.
+The nature of these costs provides a basis for distinguishing among different recommendation tasks.
+This paper aims to clarify the relationship between task formulation, solution design, and evaluation, particularly offline evaluation.
+By promoting a more structured and comprehensive understanding of these key elements, we hope to deepen the community’s appreciation of RecSys task complexity and support both new and experienced researchers in advancing the field across diverse real-world domains.
+We also urge researchers and reviewers to recognize the importance of task specificity and to value innovations tailored to distinct recommendation scenarios.